# Context-Aware Code Wiring Recommendation with LLM-based Agent

### v1

@@ V-A Threats to Validity
A threat to construct validity arises from possible bias in the manual identification of real-world code wiring instances. To address this, both the first and second authors independently inspected the data. The inter-rater agreement was high, with a Cohen’s kappa coefficient of 0.81, indicating strong consistency. Discrepancies were resolved through discussion until consensus was reached.
V-B Limitations
One limitation of WIRL is that it recommends substitutions for unresolved elements using only predefined variables available in the surrounding context. However, in some cases, the necessary variables may not exist in the current context, and declaring new variables becomes necessary. Since WIRL is specifically designed to support context-aware code adaptation, handling variable declarations falls outside its current scope. We plan to extend WIRL to address such scenarios in future work.
Another limitation is that WIRL currently supports only the Java programming language and has been evaluated solely on Java-based code. This is partly due to the limited availability of cross-language datasets involving real-world code reuse, such as from Stack Overflow to GitHub. In future research, it would be valuable to explore how to generalize and adapt WIRL to support copy-paste-modify practices in other programming languages.
@@ VI-A Code Snippet Adaptation
Numerous empirical studies and tools have been proposed to tackle code snippet adaptation. Wu et al. [48] conducted the first empirical investigation into how developers reuse Stack Overflow code, identifying five reuse patterns: exact copy, cosmetic modification, non-cosmetic modification (e.g., adding/removing statements or altering structure), idea conversion (writing from scratch), and information use. Extending this work, Chen et al. [47] examined how Stack Overflow answers are reused in GitHub projects. Using a hybrid of clone detection, keyword search, and manual inspection, they identified genuine reuse cases and proposed 11 sub-types of non-cosmetic modifications. They found that most developers reused at least some code, frequently applying statement-level or structural edits.
Zhang et al.[46] explored context-based adaptations through interviews with 21 developers, highlighting their prevalence and significance. They proposed four adaptation patterns and later introduced an interactive prompt engineering approach[30], involving executor and evaluator roles. Though iterative, this approach is not agent-based, as it cannot dynamically interact with surrounding context. Key differences from WIRL include method-level (vs. block-level) adaptation and distinct strategies for context acquisition.
Initial efforts to automate variable resolution include CSNIPPEX[29] and NLP2Testable[27], which declare default-valued variables to resolve missing references. While these address immediate issues, they lack the capacity for dynamic context adaptation, making them unsuitable for complex reuse scenarios. As Zhang et al. [46] noted, 56.1% of reuse cases require context-based adaptation, underscoring the need for more robust solutions like WIRL.
Some approaches leverage code clones and templates. Wightman et al.’s SniptMatch [24], an Eclipse plugin, maps natural language queries to snippets with variable placeholders, which users must manually fill. However, its focus is on preventing syntax errors, not enabling active adaptation. Zhang et al. [28] proposed ExampleStack, a Chrome extension that integrates online examples by leveraging past adaptations from open-source projects. It highlights frequently modified regions while preserving unchanged parts, but its dependency on existing adaptations limits its generalization to unseen cases. Lin et al. introduced CCDemon[25], which recommends edits to pasted code based on local clones, contextual elements, and naming conventions. They later proposed MICoDe[22], an Eclipse plugin for summarizing reusable module-level patterns. Similarly, EUKLAS [19] for JavaScript automatically imports missing variables and functions from the local project. While effective, these tools are constrained to local contexts and struggle with adaptation across different projects or unseen scenarios.
Another line of work treats adaptation as a deobfuscation task [20, 21], masking variables and recovering them via context-aware models pre-trained with dataflow anonymization objectives. Though conceptually aligned, these methods are computationally inefficient for scenarios involving only a few unresolved variables and are typically limited to single-file contexts, making them inadequate for broader project-level adaptation.
Other approaches support adaptation in alternative scenarios. JigSaw[26] evaluates the compatibility of inserted code via AST analysis to assist in integrating missing functionality. APIzation[23] helps convert code snippets into well-defined method declarations by statically analyzing parameters and return values. While related, these methods target different application goals than WIRL.
In contrast, WIRL enables automatic snippet adaptation with a particular emphasis on resolving unresolved variables through dynamic context analysis and LLM-powered recommendations.
VI-B LLM-based Agents for Software Engineering
LLM-based agents have become transformative tools in automating software engineering tasks through iterative reasoning and code synthesis [34, 37]. Recent work highlights their versatility: Huang et al.[31] introduced AgentCoder, a multi-agent system for code generation via dynamic testing; Lin et al.[33] modeled software processes (e.g., Scrum) by assigning LLMs to development roles (e.g., architect, tester); Zhang et al.[39] and Yang et al.[38] developed agents for test generation using code navigation toolchains. In code maintenance, Batole et al.’s LocalizeAgent[35] recommends refactorings, while Bouzenia et al.’s RepairAgent[40] performs program repair via dynamic prompt engineering. Tao et al.[36] automated GitHub issue resolution with role-specialized agents, and Ke et al.[32] applied LLM agents to flaky test repair.
@@ VI-B LLM-based Agents for Software Engineering
Despite these advancements, no prior work addresses context-aware code wiring, specifically, automatically resolving unresolved variables by substituting them with existing ones. WIRL fills this gap as the first framework to integrate agent-based planning, customized toolkit, and dynamically updated prompts for automated identifier harmonization.
@@ VII Conclusions and Future Work
In this paper, we introduced WIRL, a novel LLM-based agent designed to facilitate context-aware code wiring, an essential but underexplored aspect of the copy-paste-modify paradigm in software development. Unlike prior approaches that rely on static heuristics or historical templates, WIRL keeps a dynamically updated prompt to collect and record useful context information through a hybrid architecture that combines structured tool invocation with autonomous reasoning capabilities of LLMs. By formulating the adaptation task as an RAG infilling problem, WIRL aligns closely with the natural strengths of LLMs in code completion. Our evaluation on a carefully curated benchmark dataset demonstrates that WIRL achieves state-of-the-art performance, significantly outperforming both commercial IDEs and advanced LLMs in exact match precision and recall. These results underscore the practical value of intelligent, context-sensitive tooling for reducing developer effort and improving the reliability of reused code.
Looking ahead, we envision WIRL as a stepping stone toward more general-purpose adaptation agents that can support a broader range of integration and transformation tasks. We believe this direction holds strong potential for reshaping how developers interact with reused and AI-generated code in modern software engineering workflows.

### v2

@@ V-E Threats to Validity
The primary threat to external validity is the dataset’s size. We mitigate this by using a public, manually inspected dataset [3] and by open-sourcing our data and results [23] to encourage replication. A threat to construct validity is potential bias in identifying code wiring instances. To address this, two authors independently inspected the data, achieving a high inter-rater agreement (Cohen’s kappa of 0.81), with all discrepancies resolved through discussion.
V-F Limitations
WIRL has three main limitations. First, in the current implementation, if WIRL’s analysis does not identify any suitable existing variables with high confidence, it will not make a replacement recommendation. Instead, it will inform the developer that no suitable match was found, implicitly guiding them to declare a new variable. For future work, we plan to enhance WIRL to not only detect this scenario but also suggest contextually appropriate names and types for the new variable declaration. Second, our dataset, $CWEvaluation$, lacks more complex wiring scenarios, such as those with semantic ambiguities or many-to-many mappings, due to the laborious nature of manual collection. We plan to develop a semi-automated approach to expand the dataset in the future. Finally, WIRL is currently implemented and evaluated only for Java, and extending it to other languages remains future work.
@@ VI-A Code Snippet Adaptation
Research into code snippet adaptation began with empirical studies classifying the reuse and modification patterns developers employ when integrating external code [32, 4, 3]. Early automated tools like CSNIPPEX[33] and NLP2Testable[34] addressed unresolved variables by declaring them with default values, a solution insufficient for complex, context-dependent adaptations. A more significant body of work leverages code clones and existing examples. Tools such as SniptMatch[35], ExampleStack[16], CCDemon[36], MICoDe[37], and EUKLAS[38] recommend adaptations by matching snippets to local or online code repositories. However, their reliance on finding similar, previously adapted code limits their applicability to novel scenarios. Another strategy frames adaptation as a de-obfuscation task, masking and recovering identifiers using context-aware models[39, 40]. Other tools address related but distinct goals, such as assessing code compatibility [41] or converting snippets into methods [42].
In contrast, WIRL focuses on resolving unresolved variables in new contexts without depending on prior examples.
VI-B Code Editing
Recent advancements in code editing also inform our approach. Some models are pre-trained specifically for code modification, such as CoditT5 [43], which models edits directly, and CCT5 [44], which learns from code change datasets. Other methods leverage edit history and structural context: C3PO [45] represents edits as paths in the Abstract Syntax Tree, while Overwatch [46] learns from sequences of developer actions. More advanced LLM-based approaches like GrACE [7] and CoEdPilot [8] condition edit generation on prior associated edits to better capture developer intent.
While powerful, these methods differ from WIRL in two key ways. First, the edit location of code wiring problem is often a deterministic problem guided by compiler errors, whereas general code editing involves less predictable changes across multiple files. Second, advanced code editing approaches are dependent on a rich history of prior edits, information that is inherently scarce in the context of code wiring.
VI-C LLM-based Agents for Software Engineering
LLM-based agents are increasingly used to automate complex software engineering tasks through iterative reasoning and tool use [47, 48, 49]. They have been successfully applied to code generation and testing [50, 51, 52, 53], code maintenance and repair [54, 55, 56], and automated issue resolution [57].
@@ VI-C LLM-based Agents for Software Engineering
Despite these advancements, no prior work addresses context-aware code wiring for pasted snippets. WIRL fills this gap as the first framework to integrate agent-based planning and a customized toolkit for automated variable resolution.
@@ VII Conclusions and Future Work
We introduce WIRL, a novel LLM-based agent designed to facilitate context-aware code wiring, an essential but underexplored aspect of the copy-paste-modify paradigm in software development. Unlike prior approaches , WIRL keeps a dynamically updated prompt to collect and record useful context information through a hybrid architecture that combines structured tool invocation with autonomous reasoning capabilities of LLMs. By formulating the adaptation task as an RAG infilling problem, WIRL aligns closely with the natural strengths of LLMs in code completion. Our evaluation demonstrates that WIRL significantly outperforms commercial IDEs, advanced code editing approaches, and advanced LLMs.
Looking ahead, we envision WIRL as a stepping stone toward more general-purpose adaptation agents that can support a broader range of integration and transformation tasks.
@@ VIII Data Availability
Acknowledgment
The authors thank the reviewers for their insightful comments and constructive suggestions. This work was partially supported by the National Natural Science Foundation of China (62232003 and 62172037).